TOO SHORT CN BOND LENGTHS FOUND IN Fe, Co, AND Ni ISOCYANIDE AND CYANIDES

\underline{TSUNEO HIRANO}, REI FUKUI, and UMPEI NAGASHIMA, Grid Technology Research Center, Institute of Advanced Industrial Science and Technology, 6-9-3 Ueno, Taito-ku, Tokyo 110-0015, Japan.

We were surprised, thinking it might be a misprint, with such a short C-N bond length, 1.03(8) Å, of X̂^{6}Δ FeNC when the Lie and Dagdigian paper (2001) on FeNC by LIF^{a} came out. Ziurys and Sheridan aslo found a short CN bond length for the X̂ CoCN, 1.13133 Å, by MW, last year.^{b} For X̂^{2}Δ_{i} NiCN, Kingston and Merer^{c} has reported a little short CN bond length of 1.1591(29) Å, determined by LIF, and Sheridan and Ziurys^{d} 1.1580(8) Å by MW. Our ab initio calculated r_{e} values, though some of them are preliminary at the moment, are 1.172 (FeNC), 1.171 (CoCN), and 1.166 Å (NiCN). The discrepancy between experimental and theoretical values is in the order FeNC > CoCN > NiCN, in parallel with the expected ionic character of the metal-ligand bond and accordingly with the floppy character of the bending motion. The observed too-short CN bond lengths in these radicals probably come from the inadequate model used for the analysis of observed spectra. A new model will be proposed, and calculations of the three dimensional potential energy surfaces for X̂^{6}Δ FeNC/FeCN , which are necessary for the theoretical analysis based on the new model, are in progress at the level of MR-SDCI+Q with Roos-ANO (Fe) and aug-cc-pVQZ (C and N) basis sets.

aJ. Lie and P. J. Dagdigian, J. Chem. Phys. 114, 2137 (2001).

bP.M. Sheridan and L.M. Ziurys, University of Arizona, Private communication (2003).

cC.T. Kingston, A.J. Merer, and T.D. Varberg, J. Molc. Spectrosc., 215, 106 (2002).

dP.M. Sheridan and L.M. Ziurys, J. Chem. Phys., 118, 6370 (2003).